A die is tossed. What is the probability of tossing a number between 3 and 6?
Answer by stanbon(75887)

A die is tossed. What is the probability of tossing a number between 3 and 6?
-----------
If between means "not including" the answer is 2/6 = 1/3
--------
If between means "including" the answer is 4/6 = 2/3
=============
Cheers,
Stan H.